OpenStudy (anonymous):

x’(t) = x(t) – αy(t) y’(t) = x(t) + y(t) For which values of α do there exist straight trajectories and find the equation of the lines which contains them.

OpenStudy (anonymous):

I believe this is eigenvalue problem y(t)=x'(t)/a-x(t)/a y'(t)=x''(t)/a -x'(t)/a plug that into second equation x''(t)/a -x'(t)/a= x(t) +x'(t)/a-x(t)/a

OpenStudy (anonymous):

simplify it then solve the differential equation

OpenStudy (anonymous):

dedinitely an eigenvalue problem, but I'm used to doing it with matrices. I get an eigenvalue containing α and can't get an eigenvector unless I give α a value.

OpenStudy (experimentx):

what do you mean by linear trajectory? do you mean you are seeking the solution of type y = mx

OpenStudy (anonymous):

yeah!

OpenStudy (experimentx):

looking at this applet, my best guess is it depends on initial condition http://ocw.mit.edu/courses/mathematics/18-03sc-differential-equations-fall-2011/unit-iv-first-order-systems/qualitative-behavior-phase-portraits/LinPhasePorCursor.html and you should have real eigen vectors

OpenStudy (anonymous):

The real eigenvalues are possible for α less than three. which leaves me with quite a few options. I'm pretty sure I get two independant eigenvectors for each eigenvalue. If you plug in -1 for α you get one E-vector for each one. I'm completely lost. hahaha

OpenStudy (experimentx):

looks like more can be found here http://tutorial.math.lamar.edu/Classes/DE/PhasePlane.aspx